# Lifting the Winding Number: Precise Representation of Complex Cuts in Subspace Physics Simulations

### v1

@@ Abstract.
journal: TOGccs: Computing methodologies Physical simulation
@@ 1. Introduction
From laser-cut panels to handcrafting shapes with scissors, cuts pervade daily life. Simulating cuts in thin-walled structures such as leaves, paper or fabric requires a kinematic representation that precisely resolves time-evolving spatial discontinuities. This is a requirement that is not met by current reduced order modeling (ROM) approaches (Barbič and James, 2005).
@@
Current ROMs either lack representation or “bake in” the placement of discontinuities, therefore, they do not generalize over cut placement. For instance, mesh-based ROM mode precomputation ties the basis to the underlying discretization (Sifakis and Barbic, 2012; Fulton et al., 2019a; Shen et al., 2021). Since progressing or altering a cut changes mesh connectivity, these ROMs cannot reuse the learned basis and must again precompute offline.
@@
Indeed, efforts to represent discontinuous displacements in neural reduced simulations are nascent. Recently, Chang et al. (2023) trained on simulation snapshots of a small, manually-selected set of partial cuts, producing a displacement bases that encodes the discontinuities of partial cuts in the neural network’s weights. This limits their online reduced simulations to only the cut position seen during training, and limits the precision of the cut boundary due to the neural network’s limited capacity to precisely represent a discontinuity along a curve.
@@ 2.1. Simulating Cutting of Deformable Bodies
A common approach towards simulating cutting is refining and adapting the mesh to capture discontinuities. This approach needs to balance various considerations, such as the number of newly generated elements (Bielser and Gross, 2000; Mor and Kanade, 2000), the quality of the cut details (Busaryev et al., 2013), and material information (Chen et al., 2014). While this approach is intuitive to model, practical implementation is challenging. Remeshing can challenging to implement (Zhang et al., 2019) and computationally expensive (Obiols-Sales et al., 2023). By contrast, our approach takes a mesh-free approach through neural fields and completely bypasses the need to remesh.
@@ 2.2. Neural Methods for Simulation
Among all neural methods for physics-based simulation, our approach aligns most closely with reduced-space simulation techniques, which accelerate simulations by finding a reduced basis. These bases can either be mesh-dependent (Fulton et al., 2019b; Shen et al., 2021) or mesh-independent (Chang et al., 2023), and can be data-based (derived from simulation sequences) (Chen et al., 2023b; Zong et al., 2023) or data-free (derived from domain geometry) (Sharp et al., 2023; Modi et al., 2024).
@@ 2.3. Inside/Outside Descriptors
The winding number is a fundamental concept in mathematics, particularly in complex analysis, geometry, and topology. It measures how many times a closed curve winds around a given point in the plane, by integrating the “angle of projection” along the curve with relation to the query point. As a field over all points on the plane it is piecewise constant with jump conditions across the curve (Reinhart, 1960; Shimrat, 1962), making it a natural candidate for “strict” insideness tests. Jacobson et al. (2013) proposed its generalization (GWN) to open curves, where is harmonic with jump conditions across the curve, thereby serving as a “soft” insideness test. The GWN has been implemented for point clouds (Barill et al., 2018) and curved surfaces (Feng et al., 2023), and applied to garment modeling (Hu et al., 2018; Chi and Song, 2021), geometry processing (Xu et al., 2023; Zhou et al., 2016), and computer vision (Müller et al., 2021). Our work leverages the GWN to represent discontinuities.
@@ 3. How to represent a discontinuous field using a continuous neural network
[…]
@@
[…]
@@
[…]
@@ 4.1. Precomputation: Training
In addition to the previous data-driven training setting, our method neural field can also be trained in a data-free fashion. Following Modi et al. (2024), we trained the neural network by minimizing the elastic energy
@@ 4.2. Dynamic Subspace Simulation
where $h$ is the time step size, and $\bm{z}_{j+1}^{\text{pred}}=2\bm{z}_{j}-\bm{z}_{j-1}$. Evaluating the elastic energy $\Psi$ involves computing the deformation gradient $\mathbf{F}=\nicefrac{{\partial\bm{u}(\bm{g},\bm{x},\bm{z})}}{{\partial\bm{x}}}$. This gradient can be directly obtained from the reduced-space coordinate $\bm{z}$ and the spatial gradient of the neural field, $\nicefrac{{\partial\mathbf{B}_{\bm{x},\bm{g}}}}{{\partial\bm{x}}}$, which is efficiently calculated via automatic differentiation of the neural network. We evaluate the domain integral using uniform stochastic cubature. Our examples optimize (10) using gradient descent; we also implemented Newton’s method and observed similar performance.
@@ 4.3. Implementation
Since $\bm{\Phi}$ is a $k$-dimensional basis, we must repeat the lifting construction $k$ times. Conceptually, the simplest approach is to train $k$ volumetric networks, however, this is not strictly necessarily, since a volumetric neural field is not limited to a 1-dimensional (real-valued) output. In our implementation, we parameterize all $k$ displacement bases using the same neural network weights $\theta$ by making $\tilde{\bm{\phi}}_{\theta}^{\alpha}$ a $k$-valued field, i.e., $\bm{\phi}_{i}^{\alpha}=\tilde{\bm{\phi}}_{\theta,i}^{\alpha}\circ\mathcal{L}^{\alpha}$.
@@ 5. Results
Table 1 summarizes the timing statistics for the examples in our paper. Time per step increases with cut changes due to the need to re-evaluate the neural network and update the basis for the new cut shape. Without cut changes, our method matches traditional reduced space simulations. Our method runs at $28.5\sim 42.4$ fps with cut changes and $62\sim 166$ fps without. For reference, full-space simulation takes around 1.1s per frame without cut changes during simulation. This gives our method a 30183×30$\sim$ 183$\times$30 ∼ 183 × speedup over full-space simulations while enabling cut changes during simulation. All results are reported using an NVIDIA RTX 4090 GPU and an AMD 7950X CPU.
@@
Our method does not assume a specific type of domain discretization, instead relying only on the ability to sample the domain. All examples below use stochastic cubature drawn from a uniform distribution over the domain (Chang et al., 2023; Modi et al., 2024). Figures 10 and 13 are rendered with a mesh bind to the shape, while other figures are visualized as point clouds with colors sampled from a texture.
@@ 5.1. Evaluation on Data-Free Settings
To the best of our knowledge, our method is the first data-free model reduction method to support displacement field discontinuities, progressing cutting, or generalization of cut placement. As depicted by the butterfly in Figure 3 and the small ginkgo leaves in Figure 1, our method enables a ROM with highly detailed cut structures. All of our data-free examples use the $18$-dimensional basis presented above.
@@
To evaluate generalization over cut geometry, we trained the network weights $\theta$ on three “training” cuts, then, freezing the weights, simulated three significantly different “test” cuts, observing plausible simulated results (see Figure 7 and accompanying video).
@@ 5.2. Evaluation on Data-Driven Settings
We demonstrate generalizability with an interactive interface that enables users to edit the cut geometry while the simulation is active (see Figures 11 and 4). Starting from a straight cut, users can drag and adjust the control points of the kirigami cuts. After editing, the resulting zig-zag cut shape significantly differs from the initial straight cut, which was not included in the training data (see Figure 4).
@@
To evaluate quantitatively the generalizability of our method, we compare it with previous approaches that also use neural fields to construct the basis for reduced-space simulation, including LiCROM (Chang et al., 2023) and Simplicits (Modi et al., 2024). As a third baseline, the DANN network proposed by Belhe et al. (2023), which is designed to better capture discontinuities, replaces the standard MLP in LiCROM, keeping the remainder of LiCROM the intact.
@@
As shown in the top row of Figure 13, when evaluated on the training set, the reconstruction errors for all data-driven methods are below $1.6\times 10^{-3}$. Our approach achieves the most accurate reconstruction of deformation details and exhibits the lowest reconstruction error among them.
@@
Since Simplicits Modi et al. (2024) does not require training data, it is expected to have a larger reconstruction error. Simplicits does not consider cuts or discontinuities in its vibration mode analysis, and it also cannot leverage training data that exemplifies the displacements induced by discontinuities. For these reasons, a Simplicits model does not “know” that a cut exists. Since the neural field used in Simplicits has built-in continuity, the reduced simulations keep the two sides of the cut glued together. This result reflects the state of the art in geometry-agnostic data-free model reduction until now. By contrast, our approach is explicitly discontinuity-aware, and therefore accurately capture cutting.
When tested on cut shapes not included in training data, our method demonstrates greater generalizability. We have shown the one-shot generalizability of our approach. As shown in the bottom line of Figure 13, when tested on cut shapes with altered cut positions, all previous methods exhibit limitations in reconstructing visually reasonable results. In contrast, our method is flexible enough to adapt to new cut positions by simply updating the $\Gamma$ to the new cut position without retraining the neural basis. This adaptation is nontrivial for previous methods because their basis functions are either fully coupled to the neural network weights (Chang et al., 2023) or require remeshing and retraining of the entire domain to handle new cuts (Belhe et al., 2023).
@@ 6. Discussions and Future Work
The cut representation in this paper is limited to piecewise linear curves, which constrains its ability to model more intricate geometries. Future work could explore extending this approach to incorporate higher-order parametric representations, such as Bézier curves, as outlined in the work by Spainhour et al. (2024).

### v2

# Lifting the Winding Number: Precise Discontinuities in Neural Fields for Physics Simulation

@@ Abstract.
submissionid: 172journal: TOGjournalyear: 2025copyright: acmlicensedconference: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ; August 10–14, 2025; Vancouver, BC, Canadabooktitle: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers (SIGGRAPH Conference Papers ’25), August 10–14, 2025, Vancouver, BC, Canadadoi: 10.1145/3721238.3730597isbn: 979-8-4007-1540-2/2025/08ccs: Computing methodologies Physical simulation
@@ 1. Introduction
From laser-cut panels to handcrafting shapes with scissors, cuts pervade daily life. Simulating cuts in thin-walled structures such as leaves, paper, or fabric requires a kinematic representation that precisely resolves time-evolving spatial discontinuities. This is a requirement that is not met by current reduced order modeling (ROM) approaches (Barbič and James, 2005).
@@
Current ROMs either lack representation or “bake in” the placement of discontinuities, therefore, they do not generalize over cut placement. For instance, mesh-based ROM mode precomputation ties the basis to the underlying discretization (Sifakis and Barbic, 2012; Fulton et al., 2019; Shen et al., 2021). Since progressing or altering a cut changes mesh connectivity, these ROMs cannot reuse the learned basis and must again precompute offline.
@@
Indeed, efforts to represent discontinuous displacements in neural reduced simulations are nascent. Recently, Chang et al. (2023) trained on simulation snapshots of a small, manually-selected set of partial cuts, producing a displacement basis that encodes the discontinuities of partial cuts in the neural network’s weights. This limits their online reduced simulations to only the cut position seen during training, and limits the precision of the cut boundary due to the neural network’s limited capacity to precisely represent a discontinuity along a curve.
@@ 2.1. Simulating Cutting of Deformable Bodies
A common approach towards simulating cutting is refining and adapting the mesh to capture discontinuities. This approach needs to balance various considerations, such as the number of newly generated elements (Bielser and Gross, 2000; Mor and Kanade, 2000), the quality of the cut details (Busaryev et al., 2013), and material information (Chen et al., 2014). While this approach is intuitive to model, practical implementation is challenging. Remeshing can be challenging to implement (Zhang et al., 2019) and computationally expensive (Obiols-Sales et al., 2023). By contrast, our approach takes a mesh-free approach through neural fields and completely bypasses the need to remesh.
@@ 2.2. Neural Methods for Simulation
Among all neural methods for physics-based simulation, our approach aligns most closely with reduced-space simulation techniques, which accelerate simulations by finding a reduced basis. These bases can either be mesh-dependent (Fulton et al., 2019; Shen et al., 2021) or mesh-independent (Chang et al., 2023), and can be data-based (derived from simulation sequences) (Chen et al., 2023b; Zong et al., 2023) or data-free (derived from domain geometry) (Sharp et al., 2023; Modi et al., 2024).
@@ 2.3. Inside/Outside Descriptors
The winding number is a fundamental concept in mathematics, particularly in complex analysis, geometry, and topology. It measures how many times a closed curve winds around a given point in the plane, by integrating the “angle of projection” along the curve with relation to the query point. As a field over all points on the plane it is piecewise constant with jump conditions across the curve (Reinhart, 1960; Shimrat, 1962), making it a natural candidate for “strict” insideness tests. Jacobson et al. (2013) proposed its generalization (GWN) to open curves, where it is harmonic with jump conditions across the curve, thereby serving as a “soft” insideness test. The GWN has been implemented for point clouds (Barill et al., 2018) and curved surfaces (Feng et al., 2023), and applied to garment modeling (Hu et al., 2018; Chi and Song, 2021), geometry processing (Xu et al., 2023; Zhou et al., 2016), and computer vision (Müller et al., 2021). Our work leverages the GWN to represent discontinuities.
@@ 3. How to represent a discontinuous field using a continuous neural network
(4) $H(\mathbf{x})=\int_{0}^{1}\frac{\mathbf{\Gamma}^{\prime}(s)\cdot(\mathbf{\Gamma}(s)-\mathbf{x})^{\perp}}{|\mathbf{\Gamma}(s)-\mathbf{x}|^{2}}\,ds\ ,$
@@
(5) $H^{\alpha}(\mathbf{x})=\int_{0}^{\alpha}\frac{\mathbf{\Gamma}^{\prime}(s)\cdot(\mathbf{\Gamma}(s)-\mathbf{x})^{\perp}}{|\mathbf{\Gamma}(s)-\mathbf{x}|^{2}}\,ds\ .$
@@
(6) $\tilde{f}^{\alpha}_{\theta}:[0,1]\times\Omega\times\mathbb{R}\rightarrow\mathbb{R}:(\alpha,x,y,z)\xmapsto{\textrm{\scriptsize inference}}\tilde{f}^{\alpha}_{\theta}(x,y,z)\ .$
@@ 4.1. Precomputation: Training
In addition to the previous data-driven training setting, our neural field can also be trained in a data-free fashion. Following Modi et al. (2024), we trained the neural network by minimizing the elastic energy
@@ 4.2. Dynamic Subspace Simulation
where $h$ is the time step size, and $\bm{z}_{j+1}^{\text{pred}}=2\bm{z}_{j}-\bm{z}_{j-1}$. Evaluating the elastic energy $\Psi$ involves computing the deformation gradient $\mathbf{F}=\nicefrac{{\partial\bm{u}(\alpha,\bm{x},\bm{z})}}{{\partial\bm{x}}}$. This gradient can be directly obtained from the reduced-space coordinate $\bm{z}$ and the spatial gradient of the neural field, $\nicefrac{{\partial\bm{\phi}_{\bm{x}}^{\alpha}}}{{\partial\bm{x}}}$, which is efficiently calculated via automatic differentiation of the neural network. We evaluate the domain integral using uniform stochastic cubature. Our examples optimize (10) using gradient descent; we also implemented Newton’s method and observed similar performance.
@@ 4.3. Implementation
Since $\bm{\Phi}$ is a $k$-dimensional basis, we must repeat the lifting construction $k$ times. Conceptually, the simplest approach is to train $k$ volumetric networks, however, this is not strictly necessary, since a volumetric neural field is not limited to a 1-dimensional (real-valued) output. In our implementation, we parameterize all $k$ displacement bases using the same neural network weights $\theta$ by making $\tilde{\bm{\phi}}_{\theta}^{\alpha}$ a $k$-valued field, i.e., $\bm{\phi}_{i}^{\alpha}=\tilde{\bm{\phi}}_{\theta,i}^{\alpha}\circ\mathcal{L}^{\alpha}$.
@@ 5. Results
Table 1 summarizes the timing statistics for the examples in our paper. Time per step increases with cut changes due to the need to re-evaluate the neural network and update the basis for the new cut shape. Without cut changes, our method matches traditional reduced space simulations. Our method runs at $28.5\sim 42.4$ fps with cut changes and $62\sim 166$ fps without. For reference, full-space simulation takes around 1.1s per frame without cut changes during simulation. This gives our method a $30\times\sim 183\times$ speedup over full-space simulations while enabling cut changes during simulation. All results are reported using an NVIDIA RTX 4090 GPU and an AMD 7950X CPU. Unless otherwise noted, we use a 5-layer, 128-channel SIREN MLP with positional encoding up to maximum frequency of $2^{3}$. Winding numbers are scaled by a factor of 32 to ensure sufficient separation of lifted and 2D coordinates.
@@
Our method does not assume a specific type of domain discretization, instead relying only on the ability to sample the domain. All examples below use stochastic cubature drawn from a uniform distribution over the domain (Chang et al., 2023; Modi et al., 2024). Figures 10 and 14 are rendered with a mesh bound to the shape, while other figures are visualized as point clouds with colors sampled from a texture.
@@ 5.1. Evaluation on Data-Free Settings
To the best of our knowledge, our method is the first data-free model reduction method to support displacement field discontinuities, progressive cutting, or generalization of cut placement. As depicted by the butterfly in Figure 3 and the small ginkgo leaves in Figure 1, our method enables a ROM with highly detailed cut structures. All of our data-free examples use the $18$-dimensional basis presented above.
@@
To evaluate generalization over cut geometry, we trained the network weights $\theta$ on three “training” cuts, then froze the weights and simulated three significantly different ’test’ cuts, observing plausible results (see Figure 7 and accompanying video).
@@ 5.2. Evaluation on Data-Driven Settings
We demonstrate generalizability with an interactive interface that enables users to edit the cut geometry while the simulation is active (see Figures 11 and 4). Starting from a straight cut, users can drag and adjust the control points of the kirigami cuts. After editing, the resulting zig-zag cut shape significantly differs from the initial straight cut, which was not included in the training data (Figure 4).
@@
To evaluate quantitatively the generalizability of our method, we compare it with previous approaches that also use neural fields to construct the basis for reduced-space simulation, including LiCROM (Chang et al., 2023) and Simplicits (Modi et al., 2024). As a third baseline, the DANN network proposed by Belhe et al. (2023), which is designed to better capture discontinuities, replaces the standard MLP in LiCROM, keeping the remainder of LiCROM intact.
@@
As shown in the top row of Figure 14, when evaluated on the training set, the reconstruction errors for all data-driven methods are below $0.14\%$. Our approach achieves the most accurate reconstruction of deformation details and exhibits the lowest reconstruction error among them.
@@
Since Simplicits (Modi et al., 2024) does not require training data, it is expected to have a larger reconstruction error. Simplicits does not consider cuts or discontinuities in its vibration mode analysis, and it also cannot leverage training data that exemplifies the displacements induced by discontinuities. For these reasons, a Simplicits model does not “know” that a cut exists. Since the neural field used in Simplicits has built-in continuity, the reduced simulations keep the two sides of the cut glued together. This result reflects the state of the art in geometry-agnostic data-free model reduction until now. By contrast, our approach is explicitly discontinuity-aware, and therefore accurately captures cutting.
When tested on cut shapes not included in training data, our method demonstrates greater generalizability. We have shown the one-shot generalizability of our approach. As shown in the bottom line of Figure 14, when tested on cut shapes with altered cut positions, all previous methods exhibit limitations in reconstructing visually reasonable results. In contrast, our method is flexible enough to adapt to new cut positions by simply updating the $\Gamma$ to the new cut position without retraining the neural basis. This adaptation is nontrivial for previous methods because their basis functions are either fully coupled to the neural network weights (Chang et al., 2023) or require remeshing and retraining of the entire domain to handle new cuts (Belhe et al., 2023). We plot the reconstruction error as a function of the normalized gap difference between the test shapes and those used during training. As shown in Figure 12, the error increases with the percentage difference between training and test cuts. No significant visual artifacts are observed when the difference is below $15\%$.
@@ 6. Discussions and Future Work
The cut representation in this paper is limited to piecewise linear curves. Future work could explore extending this approach to incorporate higher-order parametric representations, such as Bézier curves (Spainhour et al., 2024).
@@
Acknowledgements.
We would like to thank our lab system administrator, John Hancock, and our financial officer, Xuan Dam, for their invaluable administrative support in making this research possible. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) grant RGPIN-2021-03733.